tem---++ vidyo New vidyo room: (old vidyo room: vidyo link PIN 1055)

News:

Meeting w/ RN and RL: need to estimate the amount of money we will spend on missions fbefore the end of the year.

Framework & production

Starting discussion on the transition from SL to CENTOS. No urgent need to dismiss condor for our queues.

Analysis

B physics

Alberto back at CERN. working on the journal answers. Deadline 2nd oct. Mario Galanti will not work on the analysis after the end of the month. He was developing the electron tagger, which is almost completed. Need to be included in the framework.

Heavy composite neutrinos

Really good news. Independent cadi line (EXO-20-011). Seeking green light. Pat found the language editor for the Preapproval. Documentation uptodate. Hoping to get everything approved before the end of the thesis. Matteo gave a talk at the Atlas Lepton+X meeting. the conveners: Antonio Sidoti and Stefania Sella.

VBS

No much news. Working on the interference gridpack generation w/ the FNAL group. Several issues to be solved on the code. No much chances to approve before the thesis.

WH_SS

SY: last friday discussed w/ Ugo and Alberto regarding the structure of the thesis. Not easy because the work was scattered. new eletron SF sttable, commisisioned for the next iteration of the latino production.

Same sign dileptons

no news

EPR

Jacopo 2 months.

Trigger

Detector

DT

Jacopo: worked with Antonio in Legnaro. Cabling the OBDT to the mu chamber. Acquired data w/ OBDT via GBTx. Starting today we will cabling a virtex 7 and one OBDT. Testing some fixes on the Montecassiano FW. Fabio still working on the concentrator board. Sandro: Antonio deployed the next generation of trigger and clock distribution (a BE board w/ kintex). In P5 we are using twinmux. Antonio is testing is the first prototype. SAndro: a chamber was extracted by Lorenzo and Franco (MB2) for reparation. the wires were removed and are going to be analysed in PD, ciemat, aachen.

Timing Layer

Time to define the title for the call for the AR position. It is very important to optimize timing to be able to have the position for december. Might be tricky. Massimo gave a presentation last week about the work they have done. Seno said that the student could come to the lab to work. Marta: students NOT associated w/ INFN do not exist for INFN.

Tracker/tracking

Mia: tracking POG meeting. presentation from Valerio from Pisa to fix cluster splitted due to inefficiencies. Need a person to take over the work from Iona, track selection via DNN. Some technical development to be done.

A.O.B.

Carlin: MB yesterday. Approved the moving of HLT on GPU. space for ideas. Configuration not defined yet. Discussed w/ the LHCC the HLT TDR , difficult to have it ready for 1st quarted 2021. Any other activity should be of secondo order. HLT TDR is clarlry the highest priority. LHCC endorsed the interest of CMS for PPS. Need to find extra money and institutes interested. INFN certainly interested. -- RobertoRossin - 2020-09-08
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2020-09-08 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback